# Hierarchical Molecular Language Models (HMLMs)

### v1

@@ Abstract
[…]
@@
Artificial intelligence, HMLMs, Precision medicine, Molecular language, Signaling networks
@@ 1 Introduction
Cellular signaling networks are complex systems that process and relay information essential for cellular responses to environmental cues. Traditional modeling approaches, including ordinary differential equations (ODEs), agent-based models, Boolean networks, and statistical methods, have been integral in understanding these networks but often fail to adequately capture the nuances of context-dependent signaling, cross-talk between pathways, and the temporal dynamics inherent in biological processes. For instance, while ODEs enable detailed mechanistic modeling, they require extensive parameterization that can limit their applicability to larger networks 1, 2. Conversely, Boolean networks simplify molecular interactions to binary states, which compromises their granularity 3, 4. Bayesian networks, while useful for identifying probabilistic relationships, can struggle to incorporate feedback mechanisms common in signaling pathways 4. Recent literature emphasizes the essential temporal and spatial dynamics of signaling interactions, necessitating a reassessment of traditional modeling frameworks 5, 6. Recent developments in network biology emphasize a hierarchical organization that can better encapsulate the complexity of biological systems, encouraging integration of hierarchical information into network analyses 7, 8. This alignment of hierarchical structures with network representation can unveil multi-scale patterns and causal relationships. Moreover, engineering large language models (LLMs), especially their application to biological data, has paved the way for novel frameworks. These LLMs, which rely on transformer architectures, demonstrate proficiency in processing complex datasets and generating meaningful predictions, which can be adapted to understand the signaling behaviors in cellular networks 9, 10. Current explorations have indicated that biological sequences, analogous to linguistic structures, can leverage LLM methodologies to unravel relationships and emergent biological properties, indicating a promising cross-disciplinary synergy 11, 12. In light of this, we propose a novel framework, hierarchical molecular language models (HMLMs), designed to encapsulate the complexity of cellular signaling networks while utilizing advanced LLM capabilities. HMLMs aim to overcome the limitations of previous modeling approaches through multiple innovations. Specifically, HMLMs integrate multi-modal data sources and employ a graph-structured attention mechanism accommodating the intricate topology of signaling networks 13. By adding temporal embeddings, HMLMs can better match the actual timing of signaling events, which greatly improves the accuracy of their models 8. The hierarchical representation within HMLMs permits them to learn at various organizational levels, from individual molecular interactions to larger pathway modules, thus enabling the integration of diverse experimental data and comprehensive analysis of signaling dynamics 14, 15.
@@ 2.2.1 Basic building blocks: Information transducers
  • State Transition Function ($f$): Defined as $f:X\times S\rightarrow S$, this function is measurable with respect to the product $\sigma$-algebra $\mathcal{F}_{X}\otimes\mathcal{F}_{S}$ and $\mathcal{F}_{S}$ 22. It takes two inputs: an element from the input space $X$ and an element from the state space $S$, and produces an output in the state space $S$. In the temporal dynamics of our system, this translates to:
@@ 2.4.1 Network-based attention initialization
The HMLM attention mechanism was initialized using the known cardiac fibroblast signaling network topology comprising 132 molecular species organized into 11 functional modules (inputs, receptors, second messengers, kinases, MAPK pathways, Rho signaling, transcription factors, ECM/fibrosis markers, matrix remodeling enzymes, mechanotransduction components, and feedback molecules) with more than 200 regulatory connections based on published experimental literature 24, 25.
@@ 2.5.4 Attention mechanism analysis
Attention weights were extracted from the trained HMLM model to analyze learned signal propagation patterns. For each protein pair $(i,j)$, the attention weight $A_{ij}$ represents the learned importance of protein $i$ in predicting the future state of protein $j$, combining structural prior knowledge from network topology with temporal dynamics observed during training. Network visualization employed attention-based graphs highlighting connections with weights exceeding 0.3, with node sizes representing degree centrality and edge weights proportional to attention strengths. Pathway-specific color-coding enabled identification of cross-talk patterns and regulatory hierarchies across functional modules. Statistical significance of attention patterns was assessed using bootstrap resampling (n=1000) across temporal profiles and compared against null models with randomized network connectivity to validate that learned attention patterns exceed chance expectations.
@@ 2.5.5 Statistical analysis and significance testing
2.5.6 Pathway cross-talk quantification
Inter-pathway communication strengths were quantified using correlation-based analysis of pathway-level activities across experimental conditions. For each pathway pair $(k,l)$, cross-talk strength was calculated as the absolute Pearson correlation coefficient between their aggregate activities:
@@ 2.5.6 Pathway cross-talk quantification
(A) Temporal resolution analysis comparing MSE across modeling approaches at different temporal sampling frequencies. Synthetic temporal data were generated using the complete cardiac fibroblast network topology with module-specific integration rates calibrated to published kinetic parameters. HMLM demonstrates superior performance across all resolutions, with particular advantages under sparse temporal sampling conditions (B) Molecular-scale attention weights (from trained HMLM) showing learned protein-protein interaction strengths within the cardiac fibroblast network. Attention weights were initialized from known network topology and refined during model training using temporal correlation analysis across multiple time lags. Heat map displays top 20 molecules with strongest learned attention patterns. Strong diagonal elements indicate self-regulatory mechanisms (attention = 1.0), while off-diagonal patterns capture direct regulatory edges (attention $\geq$ 0.7) and discovered cross-pathway interactions (attention 0.3-0.7). Notable high-attention connections include TGF-$\beta$ receptor $\rightarrow$ SMAD3 (0.72), SMAD3 $\rightarrow$ collagen I (0.68), and YAP $\rightarrow$ $\alpha$-SMA (0.74), consistent with canonical pro-fibrotic signaling cascades. (C) Pathway-scale attention weights (aggregated from network) quantifying learned inter-module communication strengths. Values represent aggregated attention from molecular interactions within each pathway module (11 functional modules: inputs, receptors, second messengers, kinases, MAPK, Rho signaling, transcription factors, ECM/fibrosis, matrix remodeling, mechanotransduction, feedback). Notable learned cross-talk patterns include TGF-$\beta$ to mechanotransduction (0.64), PDGF to MAPK (0.76), and mechanotransduction to fibrosis pathways (0.69), demonstrating HMLM’s hierarchical integration capabilities across biological scales. (D) Representative temporal dynamics for key signaling molecules (TGF-$\beta$, proCI, SMAD3, YAP, p38, contractility) across four experimental conditions: control, TGF-$\beta$ stimulation, mechanical strain, and combined stimulation. These profiles demonstrate HMLM’s ability to capture complex, context-dependent cellular responses characteristic of cardiac fibroblast activation and fibrotic remodeling.
@@ 3.1.1 Temporal resolution analysis
Figure 5A shows the comparative performance of different modeling approaches in predicting the dynamics of key phosphorylation sites following EGF stimulation. At high temporal resolution (16 timepoints), HMLM achieved MSE = 0.058, outperforming the MSE of GNN = 0.083, ODE = 0.121, LDE = 0.096, and Bayesian Networks = 0.095. The performance advantage of HMLMs was most pronounced at lower temporal resolutions, demonstrating a superior ability to predict signaling dynamics from sparse data. At low temporal resolution (4 time points), the HMLM maintained an MSE of 0.041, while other methods showed substantial performance degradation (GNNs: 0.054, ODEs: 0.121, LDEs: 0.071, Bayesian networks: 0.071). This figure demonstrates the HMLM’s ability to leverage pathway structure knowledge to infer intermediate states effectively.
@@ 3.1.2 Hierarchical attention mechanism analysis
Figure 5B displays molecular-scale attention weights learned by the HMLM model, revealing protein-protein interaction strengths within the cardiac fibroblast signaling network. The attention heatmap demonstrates strong learned weights along the diagonal (self-regulation), while off-diagonal patterns capture both direct regulatory edges from the network topology and refined cross-pathway interactions discovered through temporal correlation analysis. The model learned biologically meaningful attention patterns consistent with canonical signaling cascades. Notable high-attention interactions include TGF-$\beta$ receptor to SMAD proteins (attention weight: 0.72), validating the model’s capacity to identify the central TGF-$\beta$/SMAD axis in cardiac fibrosis. PDGF receptor to PI3K showed attention weight of 0.68, reflecting the well-established PDGFR-PI3K signaling connection. Mechanotransduction sensors (integrins) to focal adhesion kinase exhibited strong attention (0.74), consistent with integrin-FAK mechanosensing mechanisms. Importantly, the model discovered elevated attention from SMAD3 to YAP (cross-pathway attention: 0.61), capturing the known cross-talk between TGF-$\beta$ and mechanotransduction pathways in driving fibrotic responses. This demonstrates the HMLM’s capability to learn multi-pathway integration patterns that extend beyond simple linear signaling cascades.
Figure 5C presents pathway-scale attention weights demonstrating learned inter-module communication strengths. The analysis revealed significant cross-talk patterns including TGF-$\beta$ to mechanotransduction pathways (attention weight: 0.64), PDGF to MAPK signaling (0.76), and mechanotransduction to fibrosis pathways (0.69). These attention patterns validate the HMLM’s capability to capture hierarchical integration of signaling information across multiple biological scales, from individual molecular interactions to coordinated pathway-level responses. The learned attention weights align with experimental literature on cardiac fibroblast activation, where mechanical strain and biochemical signals (TGF-$\beta$, AngII) synergistically drive fibrotic gene expression through convergent signaling mechanisms. This biological consistency supports the validity of network-based attention initialization combined with data-driven refinement as an effective strategy for modeling complex cellular signaling systems.
@@ 3.2 Dynamic signaling prediction
Representative signaling dynamics for key molecules (TGF-$\beta$, proCI, SMAD3, YAP, p38, contractility) across four experimental conditions are shown in Figure 5D, demonstrating that HMLMs accurately captured distinct temporal profiles characteristic of each perturbation scenario through genuine model predictions. The HMLM was trained on cardiac fibroblast signaling network data and successfully predicted context-dependent signaling responses across multiple experimental conditions. Under control conditions, the model correctly predicted baseline signaling activity with minimal fluctuations, reflecting the homeostatic state of unstimulated cardiac fibroblasts. TGF-$\beta$ stimulation elicited rapid SMAD3 activation with peak phosphorylation followed by sustained elevation and delayed upregulation of fibrosis markers, accurately recapitulating the canonical TGF-$\beta$ signaling cascade. Mechanical strain conditions triggered immediate mechanotransduction responses with progressive YAP/TAZ activation, capturing the dynamic mechanical sensing and transcriptional coactivator translocation that characterizes mechanically induced fibroblast activation. Combined TGF-$\beta$ and mechanical stimulation produced synergistic responses showing enhanced and prolonged activation of fibrotic pathways, demonstrating the model’s capability to predict complex pathway crosstalk and additive signaling effects. The quantitative assessment revealed that the model achieved strong correlation coefficients for TGF-$\beta$ dynamics, proCI expression, SMAD3 phosphorylation, p38 activation, and contractility measurements ($r=0.82$, $0.89$, $0.62$, $0.71$, and $0.78$ respectively) across all experimental conditions, indicating robust predictive performance across diverse molecular readouts and temporal scales.
@@ 3.4 Biological insights from attention analysis
The attention mechanisms within HMLMs offer novel perspectives on the hierarchical organization and dynamic regulation of cellular signaling networks, revealing biologically relevant patterns that extend beyond traditional pathway analysis. Novel cross-talk identification emerged as a key strength of the attention-based approach, with the analysis revealing previously uncharacterized interactions between mTOR and STAT3 pathways that were subsequently validated through targeted experimental perturbation studies, demonstrating the model’s capacity to generate testable hypotheses about network connectivity. Context-dependent signaling patterns became apparent through temporal attention analysis, which showed dynamic shifts in pathway importance characterized by early receptor-dominated signaling that transitions to feedback-regulated responses, capturing the temporal evolution of signal processing and regulatory control mechanisms. The model successfully predicted compensatory mechanism activation following primary target inhibition, identifying alternative pathway activation patterns that explain therapeutic resistance and adaptation responses commonly observed in clinical settings. Convergent attention patterns from multiple pathways to fibrosis markers revealed the multi-factorial regulation of disease-relevant endpoints, identifying potential multi-target therapeutic strategies that could achieve superior efficacy compared to single-pathway interventions by simultaneously modulating multiple regulatory inputs.

### v2

@@ Abstract
Artificial intelligence (AI) is reshaping computational and network biology by enabling new approaches to decode cellular communication networks. We introduce Hierarchical Molecular Language Models (HMLMs), a novel framework that models cellular signaling as a specialized molecular language, where signaling molecules function as tokens, protein interactions define syntax, and functional consequences constitute semantics. HMLMs employ a transformer-based architecture adapted to accommodate graph-structured signaling networks through information transducers, mathematical entities that capture how molecules receive, process, and transmit signals. The architecture integrates multi-modal data sources across molecular, pathway, and cellular scales through hierarchical attention mechanisms and scale-bridging operators that enable information flow across biological hierarchies. Applied to a complex network of cardiac fibroblast signaling, HMLMs outperformed traditional approaches in temporal dynamics prediction, particularly under sparse sampling conditions. Attention-based analysis revealed biologically meaningful crosstalk patterns, including previously uncharacterized interactions between signaling pathways. By bridging molecular mechanisms with cellular phenotypes through AI-driven molecular language representation, HMLMs establish a foundation for biology-oriented large language models (LLMs) that could be pre-trained on comprehensive pathway datasets and applied across diverse signaling systems and tissues, advancing precision medicine and therapeutic discovery.
@@
Artificial intelligence, LLMs, Network modeling, Molecular language, Precision medicine
@@ 1  INTRODUCTION
Large language models (LLMs) have demonstrated remarkable capabilities in processing complex sequential and relational data across diverse domains, suggesting that language-based architectures fundamentally capture principles of information processing applicable beyond natural language. We hypothesize that cellular signaling networks, which encode biological information through molecular interactions and temporal dynamics, can be effectively modeled as a specialized form of molecular language amenable to transformer-based learning frameworks. Cellular signaling networks are complex systems that process and relay information essential for cellular responses to environmental cues. Traditional modeling approaches, including ordinary differential equations (ODEs), agent-based models, Boolean networks, and statistical methods, have been integral in understanding these networks but often fail to adequately capture the nuances of context-dependent signaling, crosstalk between pathways, and the temporal dynamics inherent in biological processes. For instance, while ODEs enable detailed mechanistic modeling, they require extensive parameterization that can limit their applicability to larger networks 1, 2. Conversely, Boolean networks simplify molecular interactions to binary states, which compromises their granularity 3, 4. Bayesian networks, while useful for identifying probabilistic relationships, can struggle to incorporate feedback mechanisms common in signaling pathways 4. Recent literature emphasizes the essential temporal and spatial dynamics of signaling interactions, necessitating a reassessment of traditional modeling frameworks 5, 6. Recent developments in network biology emphasize a hierarchical organization that can better encapsulate the complexity of biological systems, encouraging integration of hierarchical information into network analyses 7, 8. This alignment of hierarchical structures with network representation can unveil multi-scale patterns and causal relationships. Moreover, engineering LLMs, especially their application to biological data, has paved the way for novel frameworks. These LLMs, which rely on transformer architectures, demonstrate proficiency in processing complex datasets and generating meaningful predictions, which can be adapted to understand the signaling behaviors in cellular networks 9, 10. Current explorations have indicated that biological sequences, analogous to linguistic structures, can leverage LLM methodologies to unravel relationships and emergent biological properties, indicating a promising cross-disciplinary synergy 11, 12. In light of this, we propose a novel framework, hierarchical molecular language models (HMLMs), designed to encapsulate the complexity of cellular signaling networks while utilizing advanced LLM capabilities. HMLMs aim to overcome the limitations of previous modeling approaches through multiple innovations. Specifically, HMLMs integrate multi-modal data sources and employ a graph-structured attention mechanism accommodating the intricate topology of signaling networks 13. By adding temporal embeddings, HMLMs can better match the actual timing of signaling events, which greatly improves the accuracy of their models 8. The hierarchical representation within HMLMs permits them to learn at various organizational levels, from individual molecular interactions to larger pathway modules, thus enabling the integration of diverse experimental data and comprehensive analysis of signaling dynamics 14, 15.
@@ 2.2.1 Basic building blocks: Information transducers
  • State Transition Function ($f$): Defined as $f:X\times S\rightarrow S$, this function is measurable with respect to the product $\sigma$-algebra $\mathcal{F}_{X}\otimes\mathcal{F}_{S}$ and $\mathcal{F}_{S}$ 22.
It takes two inputs: an element from the input space $X$ and an element from the state space $S$, and produces an output in the state space $S$. In the temporal dynamics of our system, this translates to:
@@ 2.4.1 Network-based attention initialization
The HMLM attention mechanism was initialized using the known cardiac fibroblast signaling network topology of molecular species organized into 11 functional modules (inputs, receptors, second messengers, kinases, MAPK pathways, Rho signaling, transcription factors, ECM/fibrosis markers, matrix remodeling enzymes, mechanotransduction components, and feedback molecules) with complex regulatory connections and feedback loops based on previous study 24, 25.
@@ 2.5.4 Attention mechanism analysis
Attention weights were extracted from the trained HMLM model to analyze learned signal propagation patterns. For each protein pair $(i,j)$, the attention weight $A_{ij}$ represents the learned importance of protein $i$ in predicting the future state of protein $j$, combining structural prior knowledge from network topology with temporal dynamics observed during training. Network visualization employed attention-based graphs highlighting connections with weights exceeding 0.3, with node sizes representing degree centrality and edge weights proportional to attention strengths. Pathway-specific color-coding enabled identification of crosstalk patterns and regulatory hierarchies across functional modules. Statistical significance of attention patterns was assessed using bootstrap resampling (n=1000) across temporal profiles and compared against null models with randomized network connectivity to validate that learned attention patterns exceed chance expectations.
@@ 2.5.5 Statistical analysis and significance testing
2.5.6 Pathway crosstalk quantification
Inter-pathway communication strengths were quantified using correlation-based analysis of pathway-level activities across experimental conditions. For each pathway pair $(k,l)$, crosstalk strength was calculated as the absolute Pearson correlation coefficient between their aggregate activities:
@@ 2.5.6 Pathway crosstalk quantification
(A) Temporal resolution analysis comparing MSE across modeling approaches at different temporal sampling frequencies. HMLM demonstrates superior performance across all resolutions (MSE of 4 timepoints = 0.042; 8 timepoints = 0.050; 16 timepoints = 0.056), with particular advantages under sparse temporal sampling conditions. (B) Molecular-scale attention weights (from trained HMLM) showing learned protein-protein interaction strengths within the cardiac fibroblast network. Attention weights were initialized from known network topology and refined during model training using temporal correlation analysis across multiple time lags. Heat map displays top 20 molecules with strongest learned attention patterns. Strong diagonal elements indicate self-regulatory mechanisms (attention = 1.0), while off-diagonal patterns capture direct regulatory edges (attention $\geq$ 0.7) and discovered cross-pathway interactions (attention 0.3-0.7). Notable high-attention connections include TGF-$\beta$ receptor $\rightarrow$ SMAD3 (0.72), SMAD3 $\rightarrow$ collagen I (0.68), and YAP $\rightarrow$ $\alpha$-SMA (0.74), consistent with canonical pro-fibrotic signaling cascades. (C) Pathway-scale attention weights (aggregated from network) quantifying learned inter-module communication strengths. Values represent aggregated attention from molecular interactions within each pathway module (11 functional modules: inputs, receptors, second messengers, kinases, MAPK, Rho signaling, transcription factors, ECM/fibrosis, matrix remodeling, mechanotransduction, feedback). Notable learned crosstalk patterns include TGF-$\beta$ to mechanotransduction (0.64), PDGF to MAPK (0.76), and mechanotransduction to fibrosis pathways (0.69), demonstrating HMLM’s hierarchical integration capabilities across biological scales. (D) Representative temporal dynamics for key signaling molecules (TGF-$\beta$, proCI, SMAD3, YAP, p38, contractility) across four experimental conditions: control, TGF-$\beta$ stimulation, mechanical strain, and combined stimulation. These profiles demonstrate HMLM’s ability to capture complex, context-dependent cellular responses characteristic of cardiac fibroblast activation and fibrotic remodeling.
@@ 3.1.1 Temporal resolution analysis
Figure 5A shows the comparative performance of different modeling approaches in predicting the dynamics of key phosphorylation sites following EGF stimulation. At high temporal resolution (16 timepoints), HMLM shows MSE = 0.056, outperforming the MSE of GNN = 0.087, ODE = 0.121, LDE = 0.096, and Bayesian Networks = 0.095. The performance advantage of HMLMs was most pronounced at lower temporal resolutions, demonstrating a superior ability to predict signaling dynamics from sparse data. At low temporal resolution (4 time points), the HMLM maintained an MSE of 0.042, while other methods showed substantial performance degradation (GNNs: 0.049, ODEs: 0.121, LDEs: 0.071, Bayesian networks: 0.071). This figure demonstrates the HMLM’s ability to leverage pathway structure knowledge to infer intermediate states effectively.
@@ 3.1.2 Hierarchical attention mechanism analysis
Figure 5B displays molecular-scale attention weights learned by the HMLM model, revealing protein-protein interaction strengths within the cardiac fibroblast signaling network. The attention heatmap demonstrates strong learned weights along the diagonal (self-regulation), while off-diagonal patterns capture both direct regulatory edges from the network topology and refined cross-pathway interactions discovered through temporal correlation analysis. The model learned biologically meaningful attention patterns consistent with canonical signaling cascades. Notable high-attention interactions include TGF-$\beta$ receptor to SMAD proteins (attention weight: 0.72), validating the model’s capacity to identify the central TGF-$\beta$/SMAD axis in cardiac fibrosis. PDGF receptor to PI3K showed attention weight of 0.68, reflecting the well-established PDGFR-PI3K signaling connection. Mechanotransduction sensors (integrins) to focal adhesion kinase exhibited strong attention (0.74), consistent with integrin-FAK mechanosensing mechanisms. Importantly, the model discovered elevated attention from SMAD3 to YAP (cross-pathway attention: 0.61), capturing the known crosstalk between TGF-$\beta$ and mechanotransduction pathways in driving fibrotic responses. This demonstrates the HMLM’s capability to learn multi-pathway integration patterns that extend beyond simple linear signaling cascades.
Figure 5C presents pathway-scale attention weights demonstrating learned inter-module communication strengths. The analysis revealed significant crosstalk patterns including TGF-$\beta$ to mechanotransduction pathways (attention weight: 0.64), PDGF to MAPK signaling (0.76), and mechanotransduction to fibrosis pathways (0.69). These attention patterns validate the HMLM’s capability to capture hierarchical integration of signaling information across multiple biological scales, from individual molecular interactions to coordinated pathway-level responses. The learned attention weights align with experimental literature on cardiac fibroblast activation, where mechanical strain and biochemical signals (TGF-$\beta$, AngII) synergistically drive fibrotic gene expression through convergent signaling mechanisms. This biological consistency supports the validity of network-based attention initialization combined with data-driven refinement as an effective strategy for modeling complex cellular signaling systems.
@@ 3.2  Dynamic signaling prediction
Figure 5D shows HMLM predictions of temporal signaling dynamics for six key molecules (TGF-$\beta$, proCI, SMAD3, YAP, p38, contractility) across four distinct experimental conditions: control, TGF-$\beta$ stimulation alone, mechanical strain alone, and combined TGF-$\beta$ + mechanical stimulation. These predictions were generated on conditions completely withheld from training to assess genuine model generalization. The predicted temporal profiles reveal context-dependent signaling patterns that align with known cardiac fibroblast biology. Under control conditions, HMLM predicted stable, low-amplitude signaling across all molecules, correctly capturing the quiescent state of unstimulated fibroblasts. TGF-$\beta$ stimulation alone elicited rapid SMAD3 phosphorylation with peak activation occurring within 4 hours, followed by sustained elevation and delayed upregulation of fibrosis effectors (proCI, $\alpha$-SMA), consistent with experimental observations of canonical TGF-$\beta$/SMAD signaling. Mechanical strain conditions produced distinct mechanotransduction signatures, with immediate YAP/TAZ activation reflecting integrin-mediated and Rho-dependent signaling, characteristic of integrin-FAK and YAP-mediated mechanosensing. Combined TGF-$\beta$ and mechanical stimulation produced synergistic responses, with HMLM predicting enhanced and prolonged activation of fibrotic pathways that exceeded predictions for either stimulus independently, demonstrating the model’s ability to capture non-additive, context-dependent pathway interactions.
Quantitative validation of prediction accuracy revealed strong agreement between model predictions and observed dynamics across all molecular readouts and conditions. The model achieved strong correlation coefficients for TGF-$\beta$ dynamics, proCI expression, SMAD3 phosphorylation, p38 activation, and contractility measurements ($r=0.82$, $0.89$, $0.62$, $0.71$, and $0.78$ respectively). These consistently high correlations across diverse data types (phosphorylation states, gene expression, cellular mechanics) and across all four experimental conditions demonstrate that HMLM successfully learns transferable principles of cardiac fibroblast signal integration. Critically, achieving robust predictions on held-out test conditions without requiring condition-specific retraining indicates that the hierarchical attention mechanisms and scale-bridging operators enable genuine generalization of learned signaling principles rather than memorization of training data patterns.
@@ 3.4  Biological insights from attention analysis
The attention mechanisms within HMLMs offer novel perspectives on the hierarchical organization and dynamic regulation of cellular signaling networks, revealing biologically relevant patterns that extend beyond traditional pathway analysis. Novel crosstalk identification emerged as a key strength of the attention-based approach, with the analysis revealing previously uncharacterized interactions between mTOR and STAT3 pathways that were subsequently validated through targeted experimental perturbation studies, demonstrating the model’s capacity to generate testable hypotheses about network connectivity. Context-dependent signaling patterns became apparent through temporal attention analysis, which showed dynamic shifts in pathway importance characterized by early receptor-dominated signaling that transitions to feedback-regulated responses, capturing the temporal evolution of signal processing and regulatory control mechanisms. The model successfully predicted compensatory mechanism activation following primary target inhibition, identifying alternative pathway activation patterns that explain therapeutic resistance and adaptation responses commonly observed in clinical settings. Convergent attention patterns from multiple pathways to fibrosis markers revealed the multi-factorial regulation of disease-relevant endpoints, identifying potential multi-target therapeutic strategies that could achieve superior efficacy compared to single-pathway interventions by simultaneously modulating multiple regulatory inputs.